OpenStudy (anonymous):

Solve using the substitution method . A math test is worth 100 points . Each problem is either 3 or 4 points. How many 4 point problems are there ?

OpenStudy (anonymous):

3x + 4y = 100

OpenStudy (anonymous):

Insufficient data to form a unique solution.

OpenStudy (matt101):

ktklown is correct. How many questions are on the test? Does it say anywhere a relationship between the number of each kind of question (e.g. there are 3 more 3-point questions than 4-point questions)?

OpenStudy (anonymous):

I am so sorry , I forgot to put that there are 30 problems.

OpenStudy (anonymous):

3x + 4(30-x) = 100 3x + 120 - 4x = 100 x = 20 y = 10 10 4-points problem

OpenStudy (anonymous):

Thankyou (: